In the Hugo Award winning novel Ringworld by Larry Niven, aliens build a massive ring in space for people to live on. alo is just a rip-off of Ringworld, IMO) The 186 million mile diameter rin surrounds a star for light and energy. People inhabit the inside of the ring which is spun to simulate gravity. (a) How fast would the ring have to spin in order for your measured weight to be the same as it is on Earth? HINT: Refer back to Problem Set #5, problem #4. (b) On the Earth you are also moving in a big circle as the planet spins Shouldnt we be taking centripetal force ideas into account when we do the summation of forces on, lets say, you standing on the ground? Lets see first calculate how fast you are moving as the Earth spins [Planetary data is on the Constants and Conversions Sheet.] (c) Next, calculate your centripetal acceleration because of the rotation of Earth and compare this value to the gravitational acceleration of Earth. Make sure to comment on why we dont need to take centripetal force into account.
Answer #1

a) If the velocity of the spining of ring is V then

mV2/R = mg

given R = (186*10^{6}*1609.34)/2 m

=> V = \sqrt{Rg} = \sqrt{\frac{186*10^{6}*1609.34}{2}*9.8} = 3.87*10^{5} m/sec

b) If V is the spped by which Earth is moving then

V = \frac{2\pi R_{Earth}}{T} = \frac{2*3.14* 6400*10^{3}}{24*3600} =465.2 m/s

c)

Centripetal acceleration = V2/R =465.22/(6400*103) = 0.034 m/s2